# Multiplayer Nash Preference Optimization

### v1

@@ 3.2 Multiplayer Nash Preference Optimization
Reward-aware preference optimization (RPO) (Sun et al., 2025) demonstrates how quantitative reward signals can complement qualitative preference comparisons. This approach aligns learned implicit preference models with explicit reward models that provide graded assessments of response quality. This shift allows MNPO to move beyond binary preference margins and instead minimize discrepancies between the learned reward function $r_{\pi_{\theta}}(x,y)$ and the target explicit reward model $r^{\star}(x,y)$. Concretely, instead of optimizing only the preference order between two responses, RPO defines the loss over preference pairs as:
@@
At any iteration $t$, we construct the opponent set from a mixture of recent historical policies $\{\pi_{t-j}\}_{j=0}^{t}$ ($n\leq t+1$), weighted by coefficients $\lambda_{j}$} with $\lambda_{j}\in[0,1]$ and optionally $\sum_{j}\lambda_{j}\leq 1$. The resulting time-dependent MNPO (TD-MNPO) loss is formulated as follows:
$\mathcal{L}_{\text{TD-MNPO}}^{t,\mathbb{D}}(\pi|\beta,\{\lambda_{j}\},\eta)=\mathbb{D}\left[\log\frac{\pi(y_{w}\mid x)}{\pi(y_{l}\mid x)}-\sum_{j=0}^{n-2}\lambda_{j}\log\frac{\pi_{t-j}(y_{w}\mid x)}{\pi_{t-j}(y_{l}\mid x)}\bigg\|\,\eta\delta^{\star}\right],$ (18)
@@
This unified MNPO formulation in Eq. 18 reveals that many preference optimization algorithms can be recovered as special cases by varying the number of players $n$, choice of opponents $O_{\pi}$, distance metric $\mathbb{D}$, and target reward gap $\delta^{\star}$.
For instance, DPO emerges by setting $n=2$, $O_{\pi}=\pi_{\text{ref}}$, and $\lambda_{j}=1$. Table 1 summarizes these reductions, showing how time-dependent MNPO unifies offline and online preference optimization under one principled framework. We provide a broader overview of RLHF objectives in Appendix D.
@@ 4 Experimental Setup
We implement an online RLHF framework (Dong et al., 2024) with Gemma-2-9B-it (Team et al., 2024) as the base model. Our MNPO training consists of $T=3$ iterations, where each iteration generates responses from the current policy on a fresh prompt set and updates the policy using preference feedback. To eliminate the need for costly human annotations, we employ the reward model ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) to provide preference signals. Hyperparameter optimization proves critical, as optimal configurations vary both across different base models and between iterations of the same model. Through empirical analysis, we find that maintaining $\beta$ within the range $[0.01,10]$ consistently produces strong results. Furthermore, we observe that gradually increasing $\beta$ throughout training effectively mitigates training degradation while enabling continued model improvement. Complete implementation details and hyperparameter specifications are provided in Appendix C.
@@
In addition, we compare MNPO with a group of open-source LLMs, including LLaMA-3.1-8B-it (Dubey et al., 2024), Tulu-2-DPO-70B, LLaMA-3.3-70B-it, Mixtral-8x22B-it, and Qwen3-235B-it (Yang et al., 2025), and closed-source LLMs such as Gemini-2.5-Pro, GPT-5, and Claude-Sonnet-4.
@@
MNPO 9B 57.27 52.26 7.03
@@
MNPO 73.94 33.33 75.63 91.15 80.18 70.26 73.09 71.08
@@
MNPO 82.64 44.85 3.33 61.59 48.10
@@ 7 Conclusion
We introduce Multiplayer Nash Preference Optimization, extending Nash learning from human feedback to multiplayer settings. Our framework admits well-defined Nash equilibria and unifies existing preference optimization methods as special cases. Empirically, MNPO outperforms baselines across instruction-following and reasoning benchmarks. These results validate that multiplayer formulations better capture heterogeneous human preferences and provide more robust alignment for LLMs.
@@ Appendix B Pseudo-Algorithm of MNPO
Algorithm 1 Multiplayer Nash Preference Optimization (MNPO)
@@ Appendix C Experimental Details
For MNPO, we adopt hyperparameters consistent with SimPO and INPO, using a cosine learning-rate scheduler with a peak learning rate of $5\times 10^{-7}$, a warmup ratio of 0.1, and a global batch size of 128. The optimizer is AdamW (Loshchilov & Hutter, 2017) without weight decay. We further perform a grid search for history weights at timesteps 1 and 2, selecting from $\{0,0.1,0.333,0.5,0.667,0.9\}$. In addition, we perform a grid search for $\eta$ over $\{0.1,0.01,0.0075,0.005,0.002\}$, and set $\eta=0.0075$.
@@
Appendix D Formulations of Preference Optimization Objectives
@@
Method Objective $\pi_{\text{ref}}$ $\pi_{t}$ RL Type
@@
Appendix E Mathematical Analysis
E.1 Proof of Lemma 1
@@
Because $\mathrm{Supp}(\tilde{\pi})=\mathrm{Supp}(\pi_{\mathrm{ref}})$ and $\sum_{y}\tilde{\pi}(y\,|\,x)=1$, these pairwise ratios determine $\tilde{\pi}(\cdot\,|\,x)$ uniquely—and that unique solution is exactly the normalized distribution implied by Eq. 11 and 12, i.e., $\pi^{(t+1)}$. Hence $\tilde{\pi}=\pi^{(t+1)}$, a contradiction. Therefore the minimizer is unique. ∎
E.2 Proof of Proposition 1
@@
E.3 Theoretical Analysis of Eq. 11
Here, we provide a theoretical analysis of the iterative update in Eq.  11:
@@
By standard arguments in no-regret game dynamics (Freund & Schapire, 1999), if all players update according to Eq. 19, the joint empirical distribution converges to an $n$-player Nash equilibrium. Moreover, because the update is multiplicative in form, probabilities remain strictly positive on the support of $\pi_{\mathrm{ref}}$, preventing premature collapse.
@@
Appendix F Limitations and Future Work
Like other algorithms in the RLHF paradigm, the performance of MNPO is fundamentally linked to the quality of its preference data. Two primary limitations warrant consideration for future work.
@@ Appendix F Limitations and Future Work
F.1 External Opponent Players
@@
This formulation illustrates that MNPO with external opponent players can be viewed as an extension of knowledge distillation, where the learned policy integrates information from multiple expert models while balancing divergence from the reference model. Using the same derivation technique as DPO, we can recover Eq. 20, linking preference optimization to a generalized knowledge alignment framework.
@@
Appendix G The Use of Large Language Models (LLMs)

### v2

@@ 3.2 Multiplayer Nash Preference Optimization
Reward-aware preference optimization (RPO) (Sun et al., 2025) demonstrates how quantitative reward signals can complement qualitative preference comparisons. It aligns learned implicit preference models with explicit reward models that provide graded assessments of response quality. This shift allows MNPO to move beyond binary preference margins and instead minimize discrepancies between the learned reward function $r_{\pi_{\theta}}(x,y)$ and the target reward model $r^{\star}(x,y)$. Concretely, RPO defines the loss over preference pairs as:
@@
At any iteration $t$, we construct the opponent set from a mixture of recent time-indexed historical policies $\{\pi_{t-j}\}_{j=0}^{t}$ ($n\leq t+1$), weighted by coefficients $\lambda_{j}$} with $\lambda_{j}\in[0,1]$ and optionally $\sum_{j}\lambda_{j}\leq 1$. The resulting time-dependent MNPO (TD-MNPO) loss is formulated as follows:
$\mathcal{L}_{\text{TD}}^{t,\mathbb{D}}(\pi|\beta,\{\lambda_{j}\},\eta)=\mathbb{E}_{y,y^{\prime}\sim\pi,\,y_{w},y_{l}\sim\lambda_{\mathbb{P}}\left(y,y^{\prime}\right)}\mathbb{D}\left[\log\frac{\pi(y_{w}\mid x)}{\pi(y_{l}\mid x)}-\sum_{j=0}^{n-2}\lambda_{j}\log\frac{\pi_{t-j}(y_{w}\mid x)}{\pi_{t-j}(y_{l}\mid x)}\bigg\|\,\eta\delta^{\star}\right],$ (17)
@@
This unified MNPO formulation in Eq. 17 reveals that many preference optimization algorithms can be recovered as special cases by varying the number of players $n$, choice of opponents $O_{\pi}$, distance metric $\mathbb{D}$, and target reward gap $\delta^{\star}$. For instance, DPO emerges by setting $n=2$, $O_{\pi}=\pi_{\text{ref}}$, and $\lambda_{j}=1$. Table 1 summarizes these reductions, showing how time-dependent MNPO unifies offline and online preference optimization under one principled framework. We provide a broader overview of RLHF objectives in Appendix E.
@@
3.3 Multiplayer Game with Heterogeneous Preference Oracles
Multiplayers with Heterogeneous Preference Oracles. In many real-world alignment scenarios, preference signals originate from multiple heterogeneous sources-such as annotators with different evaluation criteria or distinct reward models trained for separate dimensions of quality (e.g., helpfulness (Tan et al., 2025), safety (Dai et al., 2023), conciseness (Dumitru et al., 2025)). We therefore generalize MNPO to the heterogeneous setting by replacing the historical mixture over past policies with a mixture over opponent policies associated with distinct preference oracles.
Concretely, each player $\pi_{i}$ is paired with a distinct reward model $r_{i}(x,y)$, inducing a player-specific preference oracle $\mathbb{P}_{i}:\mathcal{X}\times\mathcal{Y}\times\{\mathcal{Y}\}^{n-1}\rightarrow[0,1]$. The objective $J_{i}\left(\pi_{i},\left\{\pi_{j}\right\}_{j\neq i}\right)$ becomes $\mathbb{E}_{x\sim D}\left[\mathbb{E}_{y^{i}\sim\pi_{i},y^{j}\sim\pi_{j}}\left[\mathbb{P}_{i}\left(y^{i}\succ\left\{y^{j}\right\}_{j\neq i}\mid x\right)\right]-\tau\operatorname{KL}\left(\pi_{i}(\cdot\mid x)\|\pi_{\text{ref }}(\cdot\mid x)\right)\right]$, which preserves the multiplayer structure while allowing each agent to learn under its own notion of preference.
Game-Theoretic Properties. When $\mathbb{P}_{i}\neq\mathbb{P}_{j}$, the resulting game is general-sum and lacks the symmetry needed for formal Nash equilibrium guarantees. In this setting, each player has its own objective $J_{i}$ with its own oracle $P_{i}$, breaking the constant-sum structure that underpins the convergence guarantees of multiplicative weights update. Consequently, the iterative framework in Eq. 10 does not have formal convergence to the Nash equilibrium in the heterogeneous case (Daskalakis et al., 2009; Hart and Mas-Colell, 2000). To quantify the quality of a policy profile, we define a player-specific duality gap: $\operatorname{DualGap}_{i}(\pi_{i})=\max_{\pi^{\prime}_{i}\in\Pi}J_{i}(\pi^{\prime}_{i},O_{\pi})-J_{i}(\pi_{i},O_{\pi})$, where $O_{\pi}=\{\pi_{j}\}_{j\neq i}$ are the fixed opponent policies. This measures player $i$’s incentive to deviate from the current strategy. We consider the policy profile $\{\pi_{i}\}_{i=1}^{n}$ to be near a stationary point when $\max_{i}\operatorname{DualGap}_{i}(\pi_{i})\leq\epsilon$, indicating that no player has a strong incentive to deviate unilaterally. Nevertheless, the algorithmic framework remains natural and principled: each policy optimizes against the current opponent distribution using its own oracle, and as we demonstrate empirically, this approach finds effective solutions in practice.
Heterogeneous MNPO. Let $\delta_{i}^{*}$ denote the target reward gap induced by reward model $r_{i}$, following RPO’s reward-aware interpretation. The heterogeneous MNPO (HT-MNPO) update for player $i$ becomes:
$\mathcal{L}_{\text{HT}}^{i,\mathbb{D}}(\pi_{i}\mid\beta,\{\lambda_{j}\},\eta)=\mathbb{E}_{y,y^{\prime}\sim\pi_{i},y_{w},y_{l}\sim\lambda_{\mathbb{P}_{i}}\left(y,y^{\prime}\right)}\mathbb{D}\left[\log\frac{\pi_{i}(y_{w}\mid x)}{\pi_{i}(y_{l}\mid x)}-\sum_{j\neq i}\lambda_{j}\log\frac{\pi_{j}(y_{w}\mid x)}{\pi_{j}(y_{l}\mid x)}\bigg\|\eta\delta_{i}^{\star}\right],$ (18)
where the mixture is over opponent policies $\{\pi_{j}\}_{j\neq i}$ and $\{\lambda_{j}\}$ are importance weights. Eq. 18 retains the equilibrium-seeking nature of MNPO, but each policy now internalizes a reward-gap signal specific to its own reward model. As a result, MNPO can align with heterogeneous or even conflicting evaluators, potentially yielding a population equilibrium that balances multiple dimensions of quality. As empirically demonstrated later, this heterogeneous formulation achieves strong empirical performance in multi-reward-model scenarios, suggesting it finds effective stationary points even without formal equilibrium guarantees. The pseudo-algorithm is illustrated in Appendix 2.
@@ 4 Experimental Setup
We implement an online RLHF framework (Dong et al., 2024) with Gemma-2-9B-it (Team et al., 2024) as the base model. Our MNPO training consists of $T=3$ iterations, where each iteration generates responses from the current policy on a fresh prompt set and updates the policy using preference feedback. To eliminate the need for costly human annotations, we employ the reward model ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) to provide preference signals for TD-MNPO. To simulate heterogeneous preference oracles, we select Skywork-Reward-V2-Llama-3.1-8B (Liu et al., 2025) and Athene-RM-8B (Frick et al., 2024) as additional reward models for HT-MNPO. Hyperparameter optimization proves critical, as optimal configurations vary both across different base models and between iterations of the same model. Through empirical analysis, we find that maintaining $\beta$ within the range $[0.01,10]$ consistently produces strong results. Furthermore, we observe that gradually increasing $\beta$ throughout training effectively mitigates training degradation while enabling continued model improvement. Complete implementation details and hyperparameter specifications are provided in Appendix C.
@@
In addition, we compare MNPO with a group of open-source LLMs, including SmolLM3-3B (Bakouch et al., 2025), LLaMA-3.1-8B-it (Dubey et al., 2024), Olmo-2-32B-Instruct (OLMo et al., 2025), Tulu-2-DPO-70B, LLaMA-3.3-70B-it, Llama-3.1-Tulu-3-70B-DPO (Lambert et al., 2025), Mixtral-8x22B-it, and Qwen3-235B-it (Yang et al., 2025), and closed-source LLMs such as Gemini-2.5-Pro, GPT-5, and Claude-Sonnet-4.
@@
TD-MNPO 9B 57.27 52.26 7.03
HT-MNPO (ArmoRM-Llama3) 9B 57.63 50.93 7.52
HT-MNPO (Skywork-Reward-V2) 9B 56.01 50.34 7.40
HT-MNPO (Athene-RM-8B) 9B 59.64 51.17 7.07
SmolLM3-3B 3B 11.81 22.12 6.38
@@
TD-MNPO 73.94 33.33 75.63 91.15 80.18 70.26 73.09 71.08
HT-MNPO (ArmoRM-Llama3) 75.05 29.80 75.60 91.23 80.29 70.38 73.48 70.83
HT-MNPO (Skywork-Reward-V2) 75.26 36.36 75.39 91.12 80.22 70.87 73.40 71.80
HT-MNPO (Athene-RM-8B) 74.42 30.81 75.40 91.18 80.44 71.36 73.32 70.99
@@
TD-MNPO 82.64 44.85 3.33 61.59 48.10
HT-MNPO (ArmoRM-Llama3) 82.64 47.79 3.33 60.98 48.68
HT-MNPO (Skywork-Reward-V2) 82.03 49.63 0 59.76 47.86
HT-MNPO (Athene-RM-8B) 82.11 47.06 0 59.15 47.08
@@ 7 Conclusion
We introduce Multiplayer Nash Preference Optimization, extending Nash learning from human feedback to multiplayer settings. Our framework admits or approximate well-defined Nash equilibria and unifies existing preference optimization methods as special cases. Empirically, MNPO outperforms baselines across instruction-following, preference-alignment, and reasoning benchmarks. These results validate that multiplayer formulations better capture heterogeneous human preferences and provide more robust alignment for LLMs.
Acknowledgment
This work was supported in part by RS-2024-00457882, National AI Research Lab Project.
@@ Appendix B Pseudo-Algorithm of MNPO
Algorithm 1 Time-dependent Multiplayer Nash Preference Optimization (TD-MNPO)
@@
Algorithm 2 Heterogeneous Multiplayer Nash Preference Optimization (HT-MNPO)
1:Require: Number of players $n$, iterations $T$, distance metric $\mathbb{D}$, importance weights $\{\lambda_{j}\}_{j=1}^{n}$, reward scaling parameter $\eta$, regularization parameter $\beta$, initial population of policies $\{\pi_{i}^{(1)}\}_{i=1}^{n}$, reference policy $\pi_{\mathrm{ref}}$, policy class $\Pi$, heterogeneous preference oracles $\{\mathbb{P}_{i}\}_{i=1}^{n}$.
2:for iteration $t=1,2,\ldots,T$ do
3:  for player $i=1,2,\ldots,n$ do
4:   Use $\pi_{i}^{(t)}$ to generate response pairs $\big\{(y_{i,k}^{1},y_{i,k}^{2})\big\}_{k=1}^{m}$ where $y_{i,k}^{1},y_{i,k}^{2}\sim\pi_{i}^{(t)}(\cdot\mid x_{k})$.
5:   Query the heterogeneous oracle $\mathbb{P}_{i}$ to obtain the dataset $D_{t}^{(i)}=\big\{(y_{i,k}^{+},y_{i,k}^{-})\big\}_{k=1}^{m}$.
6:   Update player $i$ by $\pi_{i}^{(t+1)}\leftarrow\underset{\pi\in\Pi}{\operatorname{argmin}}\;\mathcal{L}^{i,\mathbb{D}}_{\mathrm{HT\text{-}MNPO}}$
7:  end for
8:end for
9:Output: $\left\{\pi_{i}^{(T+1)}\right\}_{i=1}^{n}$.
@@ Appendix C Experimental Details
For MNPO, we adopt hyperparameters consistent with SimPO and INPO, using a cosine learning-rate scheduler with a peak learning rate of $5\times 10^{-7}$, a warmup ratio of 0.1, and a global batch size of 128. The optimizer is AdamW (Loshchilov and Hutter, 2017) without weight decay. We further perform a grid search for history weights at timesteps 1 and 2, selecting from $\{0,0.1,0.333,0.5,0.667,0.9\}$. In addition, we perform a grid search for $\eta$ over $\{0.1,0.01,0.0075,0.005,0.002\}$, and set $\eta=0.0075$.
@@
Appendix D Additional Results
Table 5: Ablation on the number of players ($n$) in TD-MNPO, where we report AlpacaEval 2.0 (length-controlled win rate, %). Increasing $n$ consistently improves alignment quality, with diminishing returns beyond n=3n{=}3.
# Players ($n$) 1 2 3 4
AlpacaEval 2.0 53.32 54.34 (+1.02) 57.27 (+3.93) 57.42 (+4.10)
Table 6: Ablation on Different Base Models. AlpacaEval 2.0 (LC) results.
Model Llama-3-8B-it INPO TD-MNPO
AlpacaEval 2.0 24.80 41.48 (+16.64) 42.94 (+18.14)
Table 7: Mean and standard deviation of model performance across three runs under different LLM judges. This illustrates both stability and relative improvements across RLHF methods.
Model GPT-4-1106-preview GPT-4.1 GPT-5-mini
SFT Model $47.83\pm 0.92$ $41.95\pm 0.13$ $49.95\pm 0.18$
DPO $48.32\pm 1.26$ $43.87\pm 0.45$ $53.98\pm 0.34$
INPO $49.30\pm 0.95$ $47.78\pm 0.14$ $56.16\pm 0.06$
MNPO $54.05\pm 1.58$ $49.21\pm 0.14$ $57.20\pm 0.09$
Table 8: Ablation of 2-player vs. 3-player HT-MNPO on AlpacaEval 2.0. 2-player scores are averaged over all pairwise judge configurations for each reward model (e.g., Armo (Skywork) and Armo (Athene) for ArmoRM-Llama3), while 3-player results are copied from the full HT-MNPO setting in Tab. 2.
Reward model 2-player HT-MNPO 3-player HT-MNPO $\Delta$
ArmoRM-Llama3 55.43 57.63 +2.20
Skywork-Reward-V2 54.25 56.01 +1.76
Athene-RM-8B 55.71 59.64 +3.93
Appendix E Formulations of Preference Optimization Objectives
@@
Method Objective $\bm{\pi_{\text{ref}}}$ $\bm{\pi_{t}}$ RL Type
@@
Appendix F Mathematical Analysis
F.1 Proof of Lemma 1
@@
Because $\mathrm{Supp}(\tilde{\pi})=\mathrm{Supp}(\pi_{\mathrm{ref}})$ and $\sum_{y}\tilde{\pi}(y\,|\,x)=1$, these pairwise ratios determine $\tilde{\pi}(\cdot\,|\,x)$ uniquely—and that unique solution is exactly the normalized distribution implied by Eq. 10 and 11, i.e., $\pi^{(t+1)}$. Hence $\tilde{\pi}=\pi^{(t+1)}$, a contradiction. Therefore the minimizer is unique. ∎
F.2 Proof of Proposition 1
@@
F.3 Theoretical Analysis of Eq. 10
Here, we provide a theoretical analysis of the iterative update in Eq.  10:
@@
By standard arguments in no-regret game dynamics (Freund and Schapire, 1999), if all players update according to Eq. 19, the joint empirical distribution converges to an $n$-player Nash equilibrium. Moreover, because the update is multiplicative in form, probabilities remain strictly positive on the support of $\pi_{\mathrm{ref}}$, preventing premature collapse.
@@
Appendix G Limitations and Future Work
Like other algorithms in the RLHF paradigm, the performance of MNPO is fundamentally linked to the quality of its preference data. Three primary limitations warrant consideration for future work.
@@ Appendix G Limitations and Future Work
Third, our theoretical analysis focuses on the homogeneous multiplayer setting where all players share the same preference oracle. This restriction is necessary to ensure the game has constant-sum structure and that multiplicative weights updates converge to a Nash equilibrium. The heterogeneous extension (HT-MNPO), while empirically effective, operates in a general-sum game where formal convergence guarantees do not apply. Future work could explore alternative equilibrium concepts (e.g., coarse correlated equilibrium) or game structures that provide theoretical grounding for heterogeneous preference optimization.
G.1 External Opponent Players
@@
This formulation illustrates that MNPO with external opponent players can be viewed as an extension of knowledge distillation (Aminian et al., 2025), where the learned policy integrates information from multiple expert models while balancing divergence from the reference model. Using the same derivation technique as DPO, we can recover Eq. 20, linking preference optimization to a generalized knowledge alignment framework.
@@
Appendix H The Use of Large Language Models (LLMs)